A Carnot engine operates between two temperature reservoirs maintained at 493.15K and 298.15K, respectively. If the desired output of the engine is 20.00 kW, find the heat transfer to the low temperature reservoir.


1
Expert's answer
2021-06-21T05:53:07-0400
"\frac{W}{Q_H}=1-\frac{T_C}{T_H}\\
\frac{20}{Q_H}=1-\frac{298.15}{493.15}\\
Q_H=50.58\ kW"